# Optimizing Sparse SYK

### v2

@@ 1 Introduction
Simulating complex quantum mechanical systems is a potential key use-case for quantum computers [20], and for this reason understanding which classes of systems are “quantumly easy” yet “classically hard” to simulate is important in determining practical applications for future quantum devices. A particularly exciting set of physically relevant systems are quantum chemistry Hamiltonians, whose physical properties are often difficult to compute using classical algorithms [13]. These Hamiltonians are composed of low-weight fermionic interactions [11] in contrast to the setting of spin Hamiltonians more commonly studied in Hamiltonian complexity theory [1, 32, 22].
One central task in computing the properties of quantum chemistry systems is finding the ground state [13]. Though this is difficult in general—even for quantum computers—there exist classical mean-field methods such as Hartree–Fock which output Gaussian state approximations to the ground states of such systems. Understanding (I) when Gaussian states do not well-approximate the true ground state energy and (II) when quantum algorithms do is thus important for understanding where one might find a quantum advantage in a practically relevant task.
It was found in the specific case of strongly-interacting fermionic Hamiltonians that both of these properties hold for typical instances [26, 4], i.e., (I) Gaussian states provably fail and (II) quantum algorithms succeed on typical instances. A popular model for the strongly-interacting regime is the Sachdev–Ye–Kitaev (SYK) model [42, 34], representing all-to-all degree-4 fermionic interactions
@@
where each $J_{I}$ is an i.i.d. standard Gaussian and $\gamma_{I}$ is the Hermitian monomial formed by the Majorana operators indexed by the set $I$. SYK is an important fermionic Hamiltonian model due to its applications to various physical phenomena, including electron correlations in material science [31] and the holographic AdS/CFT correspondence in quantum gravity [41].
Physics-based heuristics originally showed that the ground state energy of the SYK model is $\Theta(\sqrt{n})$ (e.g., [21]); however, this was only rigorously proven by the two combined works of Feng–Tian–Wei [19] and Hastings–O’Donnell [26]. Feng, Tian, and Wei showed that the SYK Hamiltonian has an expected maximum eigenvalue [^2] Equivalent to the ground state energy up to a sign as $H_{SYK}$ and $H_{SYK}$ are distributed identically. upper-bounded by $\sqrt{\ln 2}\sqrt{n}$. Hastings and O’Donnell showed the corresponding lower bound on the maximum eigenvalue of $\Omega(\sqrt{n})$ (with high probability) by constructing an efficient quantum algorithm which certifies this lower bound for specific SYK instances.
On the other hand, Haldar–Tavakol–Scaffidi [23] showed that Gaussian states can achieve only $O(1)$ energy on typical SYK Hamiltonians. Combined with the lower bound by Hastings and O’Donnell [26], this disproves the possibility of Gaussian states achieving any constant-fraction approximation of the maximal energy. This Gaussian state energy upper bound was extended in recent work by Anschuetz, Chen, Kiani, and King [4], who showed that any high-energy state of the SYK model must have high circuit complexity. This demonstrates that any classical ansatz (whose energy can be computed on a classical computer) with an efficient representation—not just Gaussian states—cannot be close to the maximum energy of the SYK model. These results suggest that simulating the SYK model is a strong candidate for showcasing a quantum advantage.
Despite this potential for quantum advantage, it remains a major engineering challenge to simulate a Hamiltonian with $\Theta(n^{4})$ degree-4 interactions on quantum hardware. To find systems which can possibly be realized by quantum computers in the near future, it is advantageous to analyze strongly-interacting Hamiltonians with sparse interactions where each particle only interacts with a few others in total. The sparse SYK model was proposed by [49] as a modified version of SYK where each interaction term is removed independently with probability 1p1-p:
@@
Determining whether one can simulate the interesting physics of strongly-interacting Hamiltonians on near-term quantum devices using sparsified models is still unclear. Recent work attempted to simulate wormhole dynamics on a quantum computer through a specific sparse SYK model generated through machine learning [30]; however, it is an open question whether this specific sparsification actually exhibited holographic behavior [35, 39]. Additionally, Herasymenko, Stroeks, Helsen, and Terhal [28] showed that a Gaussian-approximation gap does not exist when the SYK model is too sparse. They provide an efficient classical algorithm which outputs a Gaussian state with a constant-factor approximation to the sparse SYK ground state when $p=\Theta\left(1/n^{3}\right)$. This and other previous works on sparse SYK [49, 28, 29, 39] have only considered in detail the case where $p=\Theta(1/n^{3})$, i.e., when each Majorana has in expectation a constant number of interactions. These results leave open the question of what happens between the two extremes of $p=1$ and $p=\Theta(1/n^{3})$:
@@
Our results prove that the sparse SYK model retains the maximum $1/\sqrt{n}$ Gaussian approximation gap for all $p\gtrsim 1/n^{2}$, suggesting similar physical properties between SYK and sparse SYK in this sparsity range. One such example is the fact that low-energy states of the SYK model have large amounts of magic (see [16, 27] for details on fermionic magic states); this property was known for the fully dense SYK model [25], and we show in Corollary˜4.4 that this property is robust to sparsifications through $p\gtrsim 1/n^{2}$. Our work also shows that “reasonable” properties of physical Hamiltonians like sparsity do not necessarily preclude classical hardness. This suggests the existence of other sparse and physically-relevant Hamiltonians where quantum computers can have provable advantages over classical ones. Finally, our results prove that the algorithms producing constant-approximation ratio Gaussian states when $p=\Theta(1/n^{3})$ [28, 29] fail to for any slower decay of $p$ with $n$ (up to logarithmic factors).
Figure 1: We plot the maximum achievable energies on sparse SYK Hamiltonians for various classes of quantum states (up to logarithmic factors in $p$). Green represents arbitrary disorder-independent sets of $\exp(n^{2}\log n)$ states, including states of circuit complexity $\operatorname{\widetilde{\Theta}}\left(n^{2}\right)$; blue represents Gaussian states; and red represents general quantum states generated from known quantum algorithms. Solid lines represent upper bounds proved in this work, dots represent bounds shown in previous work [23, 26, 28], dashed lines denote conjectured upper bounds in this work, and shaded regions represent where the true energy achieved by the state class may lie, i.e., where there is no tight matching lower bound.
@@ 1.1 Gaussian state upper bound
We show strong upper bounds on the optimal energy Gaussian states can achieve on sparse SYK Hamiltonians which interpolates between the $\Theta(1)$ maximum achievable energy at $p=1$ [23, 26] and the $\Theta(\sqrt{n})$ maximum achievable energy at $p=\Theta(1/n^{3})$ [28, 29].
@@
This upper-bound is tight for the sparsity range $\Omega(\log n/n^{2})\leq p\leq 1$ due to the existence of Gaussian states with energy $\Omega(1)$ for all $p$ (Subsection˜4.2). We prove Subsection˜1.1 in Subsection˜4.1 by adapting a tensor network construction from [25] originally used to prove upper bounds on ansatzes based on Gaussian states for the standard SYK model. A similar method was originally introduced in [23] to prove the $O(1)$ energy upper bound for Gaussian states on SYK. By expanding the energy calculation for an arbitrary Gaussian state utilizing Wick’s theorem, we are able to represent the value as a specific tensor network contraction. Using bounds on the operator norm of Gaussian matrices, the works of [23, 25] are able to provide energy upper bounds for arbitrary Gaussian states and specific fermionic ansatzes with low magic. In our setting, the tensors are now formed with Gaussian-times-Bernoulli random variables; we apply a stronger operator norm bound for the sparse SYK setting that allows us to achieve similar results for a large range of sparsities. As a direct corollary also due to [25], our results directly give energy upper bounds on various additional ansatzes which are based on Gaussian states for sparse SYK (Appendix˜B), including fermionic states with low magic.
@@
The circuit complexity results are achieved with the following framework: (I) We show that there exists a finite set of states which forms an $\epsilon$-net of all ansatz states in terms of energy with respect to typical sparse SYK Hamiltonians. (II) We use a Lovász theta function on graphs to bound the maximum variance of any fixed quantum state with respect to the distribution of sparse SYK Hamiltonians when $p\geq\Omega(\log n/n)$. A similar argument was shown in [4] for the original SYK model. (III) With a tight enough bound on this variance, we can use standard concentration bounds to show that the supremum of the energy over all ansatz states in the net strongly concentrates around its mean, and therefore cannot be too large. Since the supremum of the net well-approximates the supremum of all ansatz states, with high probability no ansatz state exists with high energy for typical sparse SYK instances.
@@ 1.2 Quantum algorithm
This quantum algorithm is the Hastings–O’Donnell variational algorithm [26] for optimizing the SYK model, which works as follows:
@@
The next step of the algorithm is to apply a unitary transformation on our Gaussian state to get $\rho_{\theta}\triangleq e^{-\theta\zeta}\rho_{0}e^{\theta\zeta}$, where $\zeta\triangleq\sum_{m=1}^{n_{2}}\tau_{m}\sigma_{m}$. Intuitively, we can think of this unitary transformation as a “rotation” of the $\sigma$ operators towards the $\tau$ operators in the Heisenberg picture, where the magnitude of the rotation is parameterized by the evolution time $\theta$. To first-order in $\theta$ this rotation achieves a high energy, and the main technical challenge is to upper-bound the effect of higher-order terms. Hastings–O’Donnell [26] show that for some constant $\theta\leftarrow\theta^{\ast}$ this is indeed the case for standard SYK, and $\rho_{\theta^{\ast}}$ achieves $\Omega(\sqrt{n})$ energy on both the two-color model and the original SYK model.
@@
In Section˜5 we show that this algorithm retains its performance guarantees when applied to sparse SYK Hamiltonians when $p\geq\Omega(\log n/n)$. The proof of this relies on defining a sparse two-color SYK model similar to the model in [26]. From here we apply similar analysis as the original Hastings–O’Donnell algorithm, with special care to show that the Bernoulli random variables do not affect the tails of our distribution significantly when $p$ is sufficiently large.
Combining Subsection˜1.2 with Subsection˜1.1, we show a continuous regime of sparse SYK models where we have a provable separation between states with efficient classical representations and general quantum states for the optimal achievable ratio of approximation algorithms. This separation holds for all sparse SYK Hamiltonians with $p\geq C\log n/n$, extending the analogous result shown by [26] for the special case of $p=1$.
@@ 1.3 Related work
We would like to mention the concurrent, independent work of Basso–Chen–Dalzell [8] that was released while we were preparing our manuscript. They consider a Hamiltonian ensemble similar to the sparse SYK model where the $J_{I}$ Gaussian random variables in Eq.˜2 are replaced with Rademacher random variables, and instead of each term being sparsified independently, the ensemble has a fixed number of non-zero terms $m$ and is drawn uniformly from the set of all possible subsets of size $m$.
They show similar circuit complexity lower bounds for states achieving a constant approximation (also based on the works of [14, 4]) and universality results showing a $\Theta(\sqrt{n})$ maximum eigenvalue for their Hamiltonian model. However, they do not consider specific bounds for Gaussian states like we do in our work. The main contribution of [8] is a quantum Gibbs sampling-based algorithm which produces a quantum state achieving a constant approximation to the maximum eigenvalue, for all $m>Cn\log n$. While we can only show the Hastings–O’Donnell algorithm succeeds for $m\gtrsim n^{3}\log n$, the algorithm and analysis is comparatively much simpler. Additionally, previous work has given analytic evidence that optimization algorithms based on Gibbs state preparation are non-optimal in preparing near-optimal states of certain disordered systems, in that there exist other algorithms capable of optimizing to even better energies [38, 2]. For this reason, it is important to analyze a variety of algorithms capable of achieving a constant approximation ratio.
@@ 2.1 Concentration bounds
  • Binomial Distribution Chernoff Bound [24, Theorem 18.6]: If $X\sim\text{Binomial}\left(n,p\right)$ is a Binomial random variable, for all $\delta>0$,
@@ 2.2 Fermionic operators
There exists a canonical matrix representation of the Majorana operators $\gamma_{i}$. The Weyl–Brauer matrices form a representation of the Clifford algebra, and generalize the Pauli matrices to higher dimensions [10].
@@ 2.4 Gaussian states
Gaussian states are classically useful since their energy on products of Majorana operators can be efficiently computed classically using Wick’s Theorem [48]:
@@
When considering the sparse SYK Hamiltonian, we will often be interested in a net of Gaussian states in the following sense. The relation between the gate count of a parameterized quantum circuit class and a net over the class appears in the proof of [14, Lemma D.1], which we reproduce here for completeness. However this general argument is folklore, e.g., the Solovay–Kitaev theorem.

### v3

@@ 1 Introduction
Simulating complex quantum mechanical systems is a potential key use-case for quantum computers [undef], and for this reason understanding which classes of systems are “quantumly easy” yet “classically hard” to simulate is important in determining practical applications for future quantum devices. A particularly exciting set of physically relevant systems are quantum chemistry Hamiltonians, whose physical properties are often difficult to compute using classical algorithms [undefa]. These Hamiltonians are composed of low-weight fermionic interactions [undefb] in contrast to the setting of spin Hamiltonians more commonly studied in Hamiltonian complexity theory [undefc, undefd, undefe].
One central task in computing the properties of quantum chemistry systems is finding the ground state [undefa]. Though this is difficult in general—even for quantum computers—there exist classical mean-field methods such as Hartree–Fock which output Gaussian state approximations to the ground states of such systems. Understanding (I) when Gaussian states do not well-approximate the true ground state energy and (II) when quantum algorithms do is thus important for understanding where one might find a quantum advantage in a practically relevant task.
It was found in the specific case of strongly-interacting fermionic Hamiltonians that both of these properties hold for typical instances [undeff, undefg], i.e., (I) Gaussian states provably fail and (II) quantum algorithms succeed on typical instances. A popular model for the strongly-interacting regime is the Sachdev–Ye–Kitaev (SYK) model [undefh, undefi], representing all-to-all degree-4 fermionic interactions
@@
where each $J_{I}$ is an i.i.d. standard Gaussian and $\gamma_{I}$ is the Hermitian monomial formed by the Majorana operators indexed by the set $I$. SYK is an important fermionic Hamiltonian model due to its applications to various physical phenomena, including electron correlations in material science [undefj] and the holographic AdS/CFT correspondence in quantum gravity [undefk].
Physics-based heuristics originally showed that the ground state energy of the SYK model is $\Theta(\sqrt{n})$ (e.g., [undefl]); however, this was only rigorously proven by the two combined works of Feng–Tian–Wei [undefm] and Hastings–O’Donnell [undeff]. Feng, Tian, and Wei showed that the SYK Hamiltonian has an expected maximum eigenvalue [^1] Equivalent to the ground state energy up to a sign as $H_{SYK}$ and $H_{SYK}$ are distributed identically. upper-bounded by $\sqrt{\ln 2}\sqrt{n}$. Hastings and O’Donnell showed the corresponding lower bound on the maximum eigenvalue of $\Omega(\sqrt{n})$ (with high probability) by constructing an efficient quantum algorithm which certifies this lower bound for specific SYK instances.
On the other hand, Haldar–Tavakol–Scaffidi [undefn] showed that Gaussian states can achieve only $O(1)$ energy on typical SYK Hamiltonians. Combined with the lower bound by Hastings and O’Donnell [undeff], this disproves the possibility of Gaussian states achieving any constant-fraction approximation of the maximal energy. This Gaussian state energy upper bound was extended in recent work by Anschuetz, Chen, Kiani, and King [undefg], who showed that any high-energy state of the SYK model must have high circuit complexity. This demonstrates that any classical ansatz (whose energy can be computed on a classical computer) with an efficient representation—not just Gaussian states—cannot be close to the maximum energy of the SYK model. These results suggest that simulating the SYK model is a strong candidate for showcasing a quantum advantage.
Despite this potential for quantum advantage, it remains a major engineering challenge to simulate a Hamiltonian with $\Theta(n^{4})$ degree-4 interactions on quantum hardware. To find systems which can possibly be realized by quantum computers in the near future, it is advantageous to analyze strongly-interacting Hamiltonians with sparse interactions where each particle only interacts with a few others in total. The sparse SYK model was proposed by [undefo] as a modified version of SYK where each interaction term is removed independently with probability 1p1-p:
@@
Determining whether one can simulate the interesting physics of strongly-interacting Hamiltonians on near-term quantum devices using sparsified models is still unclear. Recent work attempted to simulate wormhole dynamics on a quantum computer through a specific sparse SYK model generated through machine learning [undefp]; however, it is an open question whether this specific sparsification actually exhibited holographic behavior [undefq, undefr]. Additionally, Herasymenko, Stroeks, Helsen, and Terhal [undefs] showed that a Gaussian-approximation gap does not exist when the SYK model is too sparse. They provide an efficient classical algorithm which outputs a Gaussian state with a constant-factor approximation to the sparse SYK ground state when $p=\Theta\left(1/n^{3}\right)$. This and other previous works on sparse SYK [undefo, undefs, undeft, undefr] have only considered in detail the case where $p=\Theta(1/n^{3})$, i.e., when each Majorana has in expectation a constant number of interactions. These results leave open the question of what happens between the two extremes of $p=1$ and $p=\Theta(1/n^{3})$:
@@
Our results prove that the sparse SYK model retains the maximum $1/\sqrt{n}$ Gaussian approximation gap for all $p\gtrsim 1/n^{2}$, suggesting similar physical properties between SYK and sparse SYK in this sparsity range. One such example is the fact that low-energy states of the SYK model have large amounts of magic (see [undefu, undefv] for details on fermionic magic states); this property was known for the fully dense SYK model [undefw], and we show in Corollary˜4.4 that this property is robust to sparsifications through $p\gtrsim 1/n^{2}$. Our work also shows that “reasonable” properties of physical Hamiltonians like sparsity do not necessarily preclude classical hardness. This suggests the existence of other sparse and physically-relevant Hamiltonians where quantum computers can have provable advantages over classical ones. Finally, our results prove that the algorithms producing constant-approximation ratio Gaussian states when $p=\Theta(1/n^{3})$ [undefs, undeft] fail to for any slower decay of $p$ with $n$ (up to logarithmic factors).
Figure 1: We plot the maximum achievable energies on sparse SYK Hamiltonians for various classes of quantum states (up to logarithmic factors in $p$). Green represents arbitrary disorder-independent sets of $\exp(n^{2}\log n)$ states, including states of circuit complexity $\operatorname{\widetilde{\Theta}}\left(n^{2}\right)$; blue represents Gaussian states; and red represents general quantum states generated from known quantum algorithms. Solid lines represent upper bounds proved in this work, dots represent bounds shown in previous work [undefn, undeff, undefs], dashed lines denote conjectured upper bounds in this work, and shaded regions represent where the true energy achieved by the state class may lie, i.e., where there is no tight matching lower bound.
@@ 1.1 Gaussian state upper bound
We show strong upper bounds on the optimal energy Gaussian states can achieve on sparse SYK Hamiltonians which interpolates between the $\Theta(1)$ maximum achievable energy at $p=1$ [undefn, undeff] and the $\Theta(\sqrt{n})$ maximum achievable energy at $p=\Theta(1/n^{3})$ [undefs, undeft].
@@
This upper-bound is tight for the sparsity range $\Omega(\log n/n^{2})\leq p\leq 1$ due to the existence of Gaussian states with energy $\Omega(1)$ for all $p$ (Subsection˜4.2). We prove Subsection˜1.1 in Subsection˜4.1 by adapting a tensor network construction from [undefw] originally used to prove upper bounds on ansatzes based on Gaussian states for the standard SYK model. A similar method was originally introduced in [undefn] to prove the $O(1)$ energy upper bound for Gaussian states on SYK. By expanding the energy calculation for an arbitrary Gaussian state utilizing Wick’s theorem, we are able to represent the value as a specific tensor network contraction. Using bounds on the operator norm of Gaussian matrices, the works of [undefn, undefw] are able to provide energy upper bounds for arbitrary Gaussian states and specific fermionic ansatzes with low magic. In our setting, the tensors are now formed with Gaussian-times-Bernoulli random variables; we apply a stronger operator norm bound for the sparse SYK setting that allows us to achieve similar results for a large range of sparsities. As a direct corollary also due to [undefw], our results directly give energy upper bounds on various additional ansatzes which are based on Gaussian states for sparse SYK (Appendix˜B), including fermionic states with low magic.
@@
The circuit complexity results are achieved with the following framework: (I) We show that there exists a finite set of states which forms an $\epsilon$-net of all ansatz states in terms of energy with respect to typical sparse SYK Hamiltonians. (II) We use a Lovász theta function on graphs to bound the maximum variance of any fixed quantum state with respect to the distribution of sparse SYK Hamiltonians when $p\geq\Omega(\log n/n)$. A similar argument was shown in [undefg] for the original SYK model. (III) With a tight enough bound on this variance, we can use standard concentration bounds to show that the supremum of the energy over all ansatz states in the net strongly concentrates around its mean, and therefore cannot be too large. Since the supremum of the net well-approximates the supremum of all ansatz states, with high probability no ansatz state exists with high energy for typical sparse SYK instances.
@@ 1.2 Quantum algorithm
This quantum algorithm is the Hastings–O’Donnell variational algorithm [undeff] for optimizing the SYK model, which works as follows:
@@
The next step of the algorithm is to apply a unitary transformation on our Gaussian state to get $\rho_{\theta}\triangleq e^{-\theta\zeta}\rho_{0}e^{\theta\zeta}$, where $\zeta\triangleq\sum_{m=1}^{n_{2}}\tau_{m}\sigma_{m}$. Intuitively, we can think of this unitary transformation as a “rotation” of the $\sigma$ operators towards the $\tau$ operators in the Heisenberg picture, where the magnitude of the rotation is parameterized by the evolution time $\theta$. To first-order in $\theta$ this rotation achieves a high energy, and the main technical challenge is to upper-bound the effect of higher-order terms. Hastings–O’Donnell [undeff] show that for some constant $\theta\leftarrow\theta^{\ast}$ this is indeed the case for standard SYK, and $\rho_{\theta^{\ast}}$ achieves $\Omega(\sqrt{n})$ energy on both the two-color model and the original SYK model.
@@
In Section˜5 we show that this algorithm retains its performance guarantees when applied to sparse SYK Hamiltonians when $p\geq\Omega(\log n/n)$. The proof of this relies on defining a sparse two-color SYK model similar to the model in [undeff]. From here we apply similar analysis as the original Hastings–O’Donnell algorithm, with special care to show that the Bernoulli random variables do not affect the tails of our distribution significantly when $p$ is sufficiently large.
Combining Subsection˜1.2 with Subsection˜1.1, we show a continuous regime of sparse SYK models where we have a provable separation between states with efficient classical representations and general quantum states for the optimal achievable ratio of approximation algorithms. This separation holds for all sparse SYK Hamiltonians with $p\geq C\log n/n$, extending the analogous result shown by [undeff] for the special case of $p=1$.
@@ 1.3 Related work
We would like to mention the concurrent, independent work of Basso–Chen–Dalzell [undefx] that was released while we were preparing our manuscript. They consider a Hamiltonian ensemble similar to the sparse SYK model where the $J_{I}$ Gaussian random variables in Eq.˜2 are replaced with Rademacher random variables, and instead of each term being sparsified independently, the ensemble has a fixed number of non-zero terms $m$ and is drawn uniformly from the set of all possible subsets of size $m$.
They show similar circuit complexity lower bounds for states achieving a constant approximation (also based on the works of [undefy, undefg]) and universality results showing a $\Theta(\sqrt{n})$ maximum eigenvalue for their Hamiltonian model. However, they do not consider specific bounds for Gaussian states like we do in our work. The main contribution of [undefx] is a quantum Gibbs sampling-based algorithm which produces a quantum state achieving a constant approximation to the maximum eigenvalue, for all $m>Cn\log n$. While we can only show the Hastings–O’Donnell algorithm succeeds for $m\gtrsim n^{3}\log n$, the algorithm and analysis is comparatively much simpler. Additionally, previous work has given analytic evidence that optimization algorithms based on Gibbs state preparation are non-optimal in preparing near-optimal states of certain disordered systems, in that there exist other algorithms capable of optimizing to even better energies [undefz, undefaa]. For this reason, it is important to analyze a variety of algorithms capable of achieving a constant approximation ratio.
@@ 2.1 Concentration bounds
  • Binomial Distribution Chernoff Bound [undefab, Theorem 18.6]: If $X\sim\text{Binomial}\left(n,p\right)$ is a Binomial random variable, for all $\delta>0$,
@@ 2.2 Fermionic operators
There exists a canonical matrix representation of the Majorana operators $\gamma_{i}$. The Weyl–Brauer matrices form a representation of the Clifford algebra, and generalize the Pauli matrices to higher dimensions [undefac].
@@ 2.4 Gaussian states
Gaussian states are classically useful since their energy on products of Majorana operators can be efficiently computed classically using Wick’s Theorem [undefaf]:
@@
Denote
$\lambda_{\mathrm{Gauss}}(H)\triangleq\max_{\text{Gaussian states }\psi}\braket{\psi|H|\psi},$ (19)
@@
When considering the sparse SYK Hamiltonian, we will often be interested in a net of Gaussian states in the following sense. The relation between the gate count of a parameterized quantum circuit class and a net over the class appears in the proof of [undefy, Lemma D.1], which we reproduce here for completeness. However this general argument is folklore, e.g., the Solovay–Kitaev theorem.